Where is TFC on this graph, I know AFC is between ATC and AVC. But where would TFC be?

At the profit-maximizing output, total fixed cost 35 Dollars Multiple Choice O is fgab. O sogon is ba O is efbe
Answer #1

Answer

TFC=total fixed cost =AFC*Q
AFC=ATC-AVC
TFC=(ATC-AVC)*Q

at Q=n
ATC=f and AVC=g
AFC=f-g
ATC=(f-g)*n=area gfba

so it is option 1